# Imperfect quantum networks with tailored resource states

### v1

@@ Abstract
I Introduction
Quantum communication is at the forefront of upcoming quantum technologies, representing a cutting-edge approach to secure and efficient data transmission. The overall goal in this field is to build a quantum internet Kimble (2008); Wehner et al. (2018); Azuma et al. (2021, 2023); Cacciapuoti et al. (2020a); Illiano et al. (2022); Cacciapuoti et al. (2020b) that would allow the distribution of quantum entanglement between any two or more points on the planet. Said entanglement is a crucial resource to enable applications in different areas such as quantum cryptography Gisin et al. (2002); Shor and Preskill (2000), quantum sensing Eldredge et al. (2018); Sekatski et al. (2020), and distributed quantum computing Cirac et al. (1999); Cacciapuoti et al. (2020c). In particular, the benefits of the use of multipartite entanglement is a topic of notable interest in terms of the applications that use multipartite entangled states, e.g., conference key agreement Murta et al. (2020); Hahn et al. (2020), (multipartite) quantum secret sharing Markham and Sanders (2008); Hillery et al. (1999). However, it is also relevant to assess the benefits of the use of multipartite entanglement at any part of the distribution process. For example, entanglement-based quantum networks Pirker et al. (2018); Pirker and Dür (2019); Miguel-Ramiro et al. (2023); Mor-Ruiz and Dür (2023) and measurement-based quantum computation Raussendorf and Briegel (2001); Raussendorf et al. (2003); Briegel et al. (2009) are instances that use multipartite entangled states as resource states.
Recently, in Mor-Ruiz and Dür (2023) an efficient method to describe noisy graph transformations has been found, the so-called noisy stabilizer formalism. In the framework of quantum networks, this is extremely useful since many protocols are formulated with graph states Hein et al. (2004, 2006) or employ resource states which can be described as a graph state Hahn et al. (2019, 2022); Meignant et al. (2019); Mannalath and Pathak (2023); Fischer and Towsley (2021). Furthermore, the study of the impact of noise in the different stages of these applications is relevant to simulate and study the challenges faced in experimental realizations.
In this work, we focus on entanglement-based quantum networks, where a multipartite entangled state is distributed in a quantum network with spatially separated nodes and is then locally manipulated to achieve a request in the network, e.g., bipartite communication between two nodes. These networks correspond to a top-down approach to quantum networks, as the entanglement is established before the request is known. This pre-establishment of the resource state allows for flexibility not present in bottom-up approaches to quantum networks Matsuzaki et al. (2010); Meter et al. (2011); Epping et al. (2016); Pirandola (2019); Wallnöfer et al. (2022); Kozlowski and Wehner (2019); Coelho Coutinho et al. (2022); Bugalho et al. (2023). In particular, here we use the fact that the distribution of this resource state or entanglement topology does not need to match the structure of the network to tailor the entanglement topology to a certain set of possible requests in the network.
@@ I Introduction
The paper is organized as follows. In Sec. II we provide the necessary concepts on entanglement-based quantum networks, graph states, and their manipulations, as well as the noisy stabilizer formalism and the approach we consider for the time-dependencies in the manipulation, the so-called one-shot approach. In Sec. III we discuss the setting, which includes the used noise model and the resource and target states we consider. Next, in Sec. IV, we establish two network protocols to achieve the proposed target state considering the corresponding timings and communication of classical information. Then, we analyze the protocols under the considered setting in Sec. V, which includes the optimization of the manipulation of the resource state for a symmetric scenario, the introduction of asymmetries and their impact w.r.t. noise, followed by the analysis of a tailored resource state in an asymmetric scenario. Lastly, we summarize and conclude in Sec. VI.
II Concepts
II.1 Entanglement-based quantum networks
Entanglement-Based Quantum Networks (EBQNs) Pirker et al. (2018); Pirker and Dür (2019); Miguel-Ramiro et al. (2023); Mor-Ruiz and Dür (2023) correspond to a top-down approach to quantum networks. They utilize multipartite entangled states which are locally manipulated to fulfill network tasks on demand. Three phases can be distinguished in these networks Pirker and Dür (2019),
@@ II.1 Entanglement-based quantum networks
    Static phase: After establishing the resource state, it is stored until there is a network request (e.g., a bipartite connection between two network nodes). Given that storage using quantum memories is experimentally challenging, resource states with minimal required storage are preferred Miguel-Ramiro et al. (2023).
@@
Long waiting times for the network users are avoided due to the pre-generated entanglement, which in turn requires high-quality quantum memories in the network devices for long-term storage.
Fig. 1: Schematic depiction of the different structures of an EBQN, where c.c. is short for communication channels. Note that the classical c.c., the quantum c.c. and the entanglement do not exactly match.
One central feature of EBQNs is that the structure of the communication channels does not necessarily need to match the distribution of the entanglement of the resource states. For our purposes, we distinguish the geometry and the entanglement topology of a network. The former denotes the positioning of the nodes of the network and the communication channels, both classical and quantum. The quantum ones are mostly used to generate and distribute the resource state in the dynamic phase, and thus, are most probably located between neighboring nodes. However, the classical ones are used in the adaptive phase and might not coincide with the quantum ones. Conversely, the entanglement topology refers to the entanglement structure of the qubits of the resource state. There are many possible choices for a resource state that can fulfill a given set of requests. Therefore, one can customize the resource state to optimize the performance of the network. Fig. 1 graphically represents the different structures and channels in an EBQN.
II.2 Graph states
Graph states Hein et al. (2004, 2006) have been identified as suitable candidates for the resource states, as they allow for the transformation to other entangled states via local operations and classical communication. These states form a subclass of genuinely multipartite entangled states where a state can be associated with a graph $G=(V,E)$. Qubits are represented by the set of vertices $V$, and the set of edges between vertices $E$ denotes entangling operations between two qubits, in particular, controlled-$Z$ gates. The corresponding graph state is
@@ II.2 Graph states
Graph states can be manipulated and transformed by local Clifford operations and Pauli measurements Hein et al. (2006, 2004), where the former is a set of unitary quantum operations that map stabilizer states to stabilizer states. For example, the local complementation is a local Clifford operation that transforms the entanglement structure of a graph state such that given some vertex $v$, the sub-graph of the neighborhood of $v$ is inverted. The local Pauli measurements lead to another graph state (up to local corrections that depend on the measurement basis and outcome) where the measured qubit is disconnected. Measurement in the $\sigma_{z}$ basis simply deletes all edges associated with the measured vertex, it is also referred to as vertex deletion. The measurement in the $\sigma_{y}$ basis acts as a local complementation applied to the measured qubit followed by the deletion of it. The measurement in the $\sigma_{x}$ basis acts as a local complementation on a neighboring qubit of the measured one, then a local complementation on the measured qubit, followed by the deletion of it and lastly repeating the local complementation on the neighboring qubit of the first step.
II.3 Noisy stabilizer formalism
@@
To describe and analyze noise acting on graph states and their manipulation, the Noisy Stabilizer Formalism (NSF) was introduced in Mor-Ruiz and Dür (2023). This method allows one to accurately track the full description of the noise and, for certain important noise models such as local initial Pauli noise (e.g., see Sec. III.1), scales linearly in the number of qubits of the initial state, rather than exponentially. This is possible because it independently updates the graph states and noise operators, which in turn, reveals the contributions of different sources of noise on the final state. These updates are based on the commutation relations between Pauli noise operators and manipulation operators and are referred to as update rules. Once the noise operators and the graph state are updated, their size must match. Thus, if the target state is small, one can efficiently apply the noise maps to it and retrieve the final mixed state, thereby avoiding computation with large density matrices.
@@
This is in contrast to approaches that use the efficient classical simulation of Clifford Circuits (e.g., Aaronson and Gottesman (2004); Anders and Briegel (2006); Bravyi et al. (2019); Cirq Developers (2023); Aleksandrowicz et al. (2019); Gidney (2021)) and handle noise processes in a Monte Carlo fashion by randomly inserting additional gates. This type of approach allows to sample from the output distribution; however, it does not provide a full description of the final state. Compared to other methods of tracking noise transformation under Clifford operations Janardan et al. (2016); Miller et al. (2018), the NSF method, in particular, allows tracking individual noise contributions instead of a combined error probability vector and uses a formulation that is inherently based on graph states. In this work, we use the NSF with local Pauli noise as a noise model, which while highly structured, does not correspond to a sparse error probability vector. An implementation of this formalism for graph states has been developed using Python and it is publicly available on Wallnöfer (2024).
II.4 One-shot approach
@@
This approach is analogous to the perspective found in measurement-based quantum computation Raussendorf and Briegel (2001); Raussendorf et al. (2003); Briegel et al. (2009), which also considers measurements on a graph state that acts as a resource. Despite the measurement outcomes being random, the final state or outcome of the computation can be deterministically retrieved independently of the order and outcomes of the measurements, as long as only Pauli measurements are used.
Two key insights allow this. The correction operations, which are necessary to retrieve a graph state after a measurement, are always local Clifford operators, i.e., local basis changes of the affected qubits. So, instead of performing the correction operation, one can simply change the basis of the measurement accordingly, as due to the properties of the Clifford group, any local Pauli measurement is always mapped to the same or a different local Pauli measurement. Furthermore, the local Clifford operators associated with different measurement outcomes are the same up to local Pauli operators (see e.g., Hein et al. (2006)), which can potentially swap the sign of a Pauli measurement, i.e., they can change the interpretation of the measurement outcome, but not the basis in which one measures. In Appendix A, we give an explicit computation of these measurement patterns and their equivalence with the strategies, we also analyze how this approach can be simulated using the NSF.
III Setting
@@
We consider a network consisting of $N$ nodes where a cluster state is distributed as a minimal-storage resource state. In particular, we study a single bipartite communication as the request, as it is a useful target state in quantum communication tasks, such as quantum key distribution Ekert (1991). This target state, the bipartite graph, is a Bell state up to a local basis change.
III.1 Noise model
@@
where $q_{v}(t)$ denotes the time-dependent noise probability of the storage. Following Hein et al. (2005), we define this coefficient as $q_{v}(t)=\left(1-e^{-t/T_{v}}\right)/2$, where $t$ is a certain time in the network, and $T_{v}$ is the dephasing time, which we use to describe the quality of the memories in the network. The qubits in the network are expected to be affected by this noise for different amounts of time, e.g., until being measured or while waiting for all necessary information to arrive, which naturally depends on the protocol that is used. Nevertheless, Eq. (4) is valid for any time interval $t$. Therefore, each qubit is subject to $\mathcal{E}_{v}=\mathcal{D}_{v}\mathcal{Z}_{v}$, then an $N$-qubit noisy resource state is
@@
The use of single-qubit noise models in this scenario is reasonable as we assume that during the static phase, the qubits are stored in spatially separated nodes. Moreover, one candidate method for resource generation and distribution in EBQNs is the use of entanglement purification, which leads to a nearly local noise model Wallnöfer and Dür (2017). Furthermore, local noisy measurements are accurately modeled using local depolarizing noise followed by a perfect measurement. Overall, the depolarizing noise model is considered general enough to accurately describe uncorrelated noise effects and can be understood as a worst-case model for single-qubit noise Dür et al. (2005). The use of time-dependent dephasing noise is widely used to describe the decoherence process of a quantum memory Razavi et al. (2009); Luong et al. (2016); Wallnöfer et al. (2022), and there are indications that dephasing errors are the limiting factor in diamond-based nuclear-spin memories Kalb et al. (2018); Pompili et al. (2021). Importantly, the two noise models we consider are Pauli-diagonal noise channels, such that their impact on graph states can be efficiently described using the NSF.
III.2 Entanglement topology and manipulation
@@
As mentioned in Sec. II, the choice of the resource state in EBQNs must minimize storage; therefore, we choose a state with a single qubit per node. In Mor-Ruiz and Dür (2023), several structures under depolarizing noise were studied as resource states to achieve a Bell pair with the conclusion that high-dimensional cluster states are good candidates for this task. So, we consider an $N$-qubit $d$-dimensional cluster state as a resource state.
@@
To achieve a Bell pair from a cluster state, first, we choose the path to connect the two target qubits. We aim for the shortest one as the results in Mor-Ruiz and Dür (2023) show that the number of measurements required in the manipulation directly impacts the noise of the Bell pair. There are two kinds of qubits involved in the manipulation, the ones surrounding the path and the ones inside the path. We refer to them as outer and inner neighbors, respectively. The outer neighbors are measured in the $\sigma_{z}$ basis, as one wants to disconnect the target qubits and the inner neighbors from the rest of the cluster. This leads to a one-dimensional cluster with the targets at the ends and the inner neighbors in between. Then, for the measurement of the inner neighbors, multiple sets of measurement patterns achieve the target Bell pair. The choice of such a measurement pattern leads to different noise maps acting on the target Bell pair, which in turn may result in different fidelities. Therefore, we do not consider one fixed measurement pattern, but we rather analyze the impact of choosing different measurement patterns to pick an optimal one for the particular setup. In Fig. 2, we graphically represent the manipulation for the one-dimensional case.
@@
Fig. 2: Manipulation of a 1D cluster into a Bell pair. The circles are qubits, the solid lines represent the entanglement between qubits and the dotted lines represent the extension of the cluster following the presented structure, so only part of a large cluster is explicitly drawn. At the top, the initial situation is depicted, where the target qubits (filled circles) have been chosen. Also, the corresponding manipulation is established, which consists of the measurement in the $\sigma_{z}$ basis of the outer neighbors and applying a measurement pattern to the inner neighbors. At the bottom, the manipulation has been performed and the resulting Bell pair and the remaining graph state are presented.
III.3 One-dimensional cluster as a building block

### v2

@@ Abstract
1 Introduction
Quantum communication is at the forefront of upcoming quantum technologies, representing a cutting-edge approach to secure and efficient data transmission. The overall goal in this field is to build a quantum internet [1, 2, 3, 4, 5, 6, 7] that would allow the distribution of quantum entanglement between any two or more points on the planet. Said entanglement is a crucial resource to enable applications in different areas such as quantum cryptography [8, 9], quantum sensing [10, 11], and distributed quantum computing [12, 13]. In particular, the benefits of the use of multipartite entanglement is a topic of notable interest in terms of the applications that use multipartite entangled states, e.g., conference key agreement [14, 15], (multipartite) quantum secret sharing [16, 17]. However, it is also relevant to assess the benefits of the use of multipartite entanglement at any part of the distribution process. For example, entanglement-based quantum networks [18, 19, 20, 21] and measurement-based quantum computation [22, 23, 24] are instances that use multipartite entangled states as resource states.
Recently, in [25] an efficient method to describe noisy graph transformations has been found, the so-called noisy stabilizer formalism. In the framework of quantum networks, this is extremely useful since many protocols are formulated with graph states [26, 27] or employ resource states which can be described as a graph state [28, 29, 30, 31, 32]. Furthermore, the study of the impact of noise in the different stages of these applications is relevant to simulate and study the challenges faced in experimental realizations.
In this work, we focus on entanglement-based quantum networks. The central idea for this approach is the distribution of a multipartite entangled state across spatially separated network nodes, which is then locally manipulated to achieve tasks for the network, e.g., establishing bipartite quantum communication between two nodes. This approach is motivated by the improvement of quantum memories in recent years [33, 34, 35, 36] and corresponds to a top-down approach to quantum networks, as the entanglement is established before the request is known. This pre-establishment of the resource state allows for (i) short waiting times for the network users, since the resource state only needs to be locally manipulated to fulfill a network request and (ii) flexibility not present in bottom-up approaches to quantum networks [37, 38, 39, 40, 41, 42, 43, 44]. In particular, here we use the fact that the distribution of this resource state or entanglement topology does not need to match the structure of the network to tailor the entanglement topology to a certain set of possible requests in the network.
@@ 1 Introduction
The paper is organized as follows. In Sec. 2 we provide the necessary concepts on entanglement-based quantum networks, graph states, and their manipulations, as well as the noisy stabilizer formalism and the approach we consider for the time-dependencies in the manipulation, the so-called one-shot approach. In Sec. 3 we discuss the setting, which includes the used noise model and the resource and target states we consider. Next, in Sec. 4, we establish two network protocols to achieve the proposed target state considering the corresponding timings and communication of classical information. Then, we analyze the protocols under the considered setting in Sec. 5, which includes the optimization of the manipulation of the resource state for a symmetric scenario, the introduction of asymmetries and their impact w.r.t. noise, followed by the analysis of a tailored resource state in an asymmetric scenario. Lastly, we summarize and conclude in Sec. 6.
2 Concepts
2.1 Entanglement-based quantum networks
Entanglement-Based Quantum Networks (EBQNs) [18, 19, 20, 21] correspond to a top-down approach to quantum networks. They utilize multipartite entangled states which are locally manipulated to fulfill network tasks on demand. Three phases can be distinguished in these networks [19],
@@ 2.1 Entanglement-based quantum networks
    Static phase: After establishing the resource state, it is stored until there is a network request (e.g., a bipartite connection between two network nodes). In order to minimize memory resources, states with minimal required storage are preferred [20].
@@
Long waiting times for the network users are avoided due to the pre-generated entanglement, which in turn requires high-quality quantum memories in the network devices for long-term storage. Note that this entanglement allows for flexibility in the network requests, since a resource state can be optimized to be able to fulfill many different requests.
[…]
[width=]Fig_1.pdf
Figure 1: Schematic depiction of the different structures of an EBQN, where c.c. is short for communication channels. Note that the classical c.c., the quantum c.c. and the entanglement do not exactly match.
One central feature of EBQNs is that the structure of the communication channels does not necessarily need to match the distribution of the entanglement of the resource states. For our purposes, we distinguish the geometry and the entanglement topology of a network. The former denotes the positioning of the nodes of the network and the communication channels, both classical and quantum. The quantum ones are mostly used to generate and distribute the resource state in the dynamic phase, and thus, are most probably located between neighboring nodes. However, the classical ones are used in the adaptive phase and might not coincide with the quantum ones. Conversely, the entanglement topology refers to the entanglement structure of the qubits of the resource state. There are many possible choices for a resource state that can fulfill a given set of requests. Therefore, one can customize the resource state to optimize the performance of the network. Fig. 1 graphically represents the different structures and channels in an EBQN. Note that the processes in the dynamic phase are highly dependent on hardware, and thus, can differ greatly, e.g., see [30, 32, 45, 44, 46, 47, 48] for graph, GHZ, and W states distribution.
2.2 Graph states
Graph states [26, 27] have been identified as suitable candidates for the resource states, as they allow for the transformation to other entangled states via local operations and classical communication. These states form a subclass of genuinely multipartite entangled states where a state can be associated with a graph $G=(V,E)$. Qubits are represented by the set of vertices $V$, and the set of edges between vertices $E$ denotes entangling operations between two qubits, in particular, controlled-$Z$ gates. The corresponding graph state is
@@ 2.2 Graph states
Graph states can be manipulated and transformed by local Clifford operations and Pauli measurements [27, 26], where the former is a set of unitary quantum operations that map stabilizer states to stabilizer states. For example, the local complementation is a local Clifford operation that transforms the entanglement structure of a graph state such that given some vertex $v$, the sub-graph of the neighborhood of $v$ is inverted. The local Pauli measurements lead to another graph state (up to local corrections that depend on the measurement basis and outcome) where the measured qubit is disconnected. Measurement in the $\sigma_{z}$ basis simply deletes all edges associated with the measured vertex, it is also referred to as vertex deletion. The measurement in the $\sigma_{y}$ basis acts as a local complementation applied to the measured qubit followed by the deletion of it. The measurement in the $\sigma_{x}$ basis acts as a local complementation on a neighboring qubit of the measured one, then a local complementation on the measured qubit, followed by the deletion of it and lastly repeating the local complementation on the neighboring qubit of the first step.
2.3 Noisy stabilizer formalism
@@
To describe and analyze noise acting on graph states and their manipulation, the Noisy Stabilizer Formalism (NSF) was introduced in [25]. This method allows one to accurately track the full description of Pauli-diagonal noise and, for certain important noise models such as local initial Pauli noise (e.g., see Sec. 3.1), scales linearly in the number of qubits of the initial state, rather than exponentially. This is possible because it independently updates the graph states and noise operators, which in turn, reveals the contributions of different sources of noise on the final state. These updates are based on the commutation relations between Pauli noise operators and manipulation operators and are referred to as update rules. Once the noise operators and the graph state are updated, their size must match. Thus, if the target state is small, one can efficiently apply the noise maps to it and retrieve the final mixed state, thereby avoiding computation with large density matrices. Besides the size of the target state, the number of terms in each of the individual noise maps and the vertex degree of the considered graph also affects the efficiency of the method. The vertex degree is relevant due to the fact that noise operators acting in a single qubit might spread to a noise operator acting on several qubits, potentially leading to more required noise map updates going forward.
@@
This is in contrast to approaches that use the efficient classical simulation of Clifford Circuits (e.g., [49, 50, 51, 52, 53, 54]) and handle noise processes in a Monte Carlo fashion by randomly inserting additional gates. This type of approach allows to sample from the output distribution; however, it does not provide a full description of the final state. Compared to other methods of tracking noise transformation under Clifford operations [55, 56], the NSF method, in particular, allows tracking individual noise contributions instead of a combined error probability vector and uses a formulation that is inherently based on graph states. In this work, we use the NSF with local Pauli noise as a noise model, which while highly structured, does not correspond to a sparse error probability vector. An implementation of this formalism for graph states has been developed using Python and it is publicly available on [57].
2.4 One-shot approach
@@
This approach is analogous to the perspective found in measurement-based quantum computation [22, 23, 24], which also considers measurements on a graph state that acts as a resource. Despite the measurement outcomes being random, the final state or outcome of the computation can be deterministically retrieved independently of the order and outcomes of the measurements, as long as only Pauli measurements are used.
Two key insights allow this. The correction operations, which are necessary to retrieve a graph state after a measurement, are always local Clifford operators, i.e., local basis changes of the affected qubits. So, instead of performing the correction operation, one can simply change the basis of the measurement accordingly, as due to the properties of the Clifford group, any local Pauli measurement is always mapped to the same or a different local Pauli measurement. Furthermore, the local Clifford operators associated with different measurement outcomes are the same up to local Pauli operators (see, e.g., [27]), which can potentially swap the sign of a Pauli measurement, i.e., they can change the interpretation of the measurement outcome, but not the basis in which one measures. In Appendix A, we give an explicit computation of these measurement patterns and their equivalence with the strategies, we also analyze how this approach can be simulated using the NSF.
3 Setting
@@
We consider a network consisting of $N$ nodes where a cluster state is distributed as a minimal-storage resource state. In particular, we study a single bipartite communication as the request, as it is a useful target state in quantum communication tasks, such as quantum key distribution [58]. This target state, the bipartite graph, is a Bell state up to a local basis change.
3.1 Noise model
@@
where $q_{v}(t)$ denotes the time-dependent noise probability of the storage. Following [59], we define this coefficient as $q_{v}(t)=\left(1-e^{-t/T_{v}}\right)/2$, where $t$ is a certain time in the network, and $T_{v}$ is the dephasing time, which we use to describe the quality of the memories in the network. The qubits in the network are expected to be affected by this noise for different amounts of time, e.g., until being measured or while waiting for all necessary information to arrive, which naturally depends on the protocol that is used. Nevertheless, Eq. (4) is valid for any time interval $t$. Therefore, each qubit is subject to $\mathcal{E}_{v}=\mathcal{D}_{v}\mathcal{Z}_{v}$, then an $N$-qubit noisy resource state is
@@
The use of single-qubit noise models in this scenario is reasonable as we assume that during the static phase, the qubits are stored in spatially separated nodes. Moreover, one candidate method for resource generation and distribution in EBQNs is the use of entanglement purification, which leads to a nearly local noise model [60]. Furthermore, local noisy measurements are accurately modeled using local depolarizing noise followed by a perfect measurement. Overall, the depolarizing noise model is considered general enough to accurately describe uncorrelated noise effects and can be understood as a worst-case model for single-qubit noise [61]. The use of time-dependent dephasing noise is widely used to describe the decoherence process of a quantum memory [62, 63, 41], and there are indications that dephasing errors are the limiting factor in diamond-based nuclear-spin memories [64, 65]. Importantly, the two noise models we consider are Pauli-diagonal noise channels, such that their impact on graph states can be efficiently described using the NSF.
3.2 Entanglement topology and manipulation
@@
As mentioned in Sec. 2, the choice of the resource state in EBQNs must minimize storage; therefore, we choose a state with a single qubit per node. In [21], several structures under a noise model consisting only of single-qubit depolarizing noise that ignores memory decoherence and timing and spacing effects were studied as resource states to achieve a Bell pair with the conclusion that high-dimensional cluster states are good candidates for this task. So, in this work consider an $N$-qubit $d$-dimensional cluster state as a resource state with a noise model that incorporates the same depolarizing noise and additional memory decoherence.
@@
To achieve a Bell pair from a cluster state, first, we choose the path to connect the two target qubits. We aim for the shortest one as the results in [21] show that the number of measurements required in the manipulation directly impacts the noise of the Bell pair. There are two kinds of qubits involved in the manipulation, the ones surrounding the path and the ones inside the path. We refer to them as outer and inner neighbors, respectively. The outer neighbors are measured in the $\sigma_{z}$ basis, as one wants to disconnect the target qubits and the inner neighbors from the rest of the cluster. This leads to a one-dimensional cluster with the targets at the ends and the inner neighbors in between. Then, for the measurement of the inner neighbors, multiple sets of measurement patterns achieve the target Bell pair. The choice of such a measurement pattern leads to different noise maps acting on the target Bell pair, which in turn may result in different fidelities. Therefore, we do not consider one fixed measurement pattern, but we rather analyze the impact of choosing different measurement patterns to pick an optimal one for the particular setup. In Fig. 2, we graphically represent the manipulation for the one-dimensional case.
[…]
[width=]Fig_2.pdf
@@
Figure 2: Manipulation of a 1D cluster into a Bell pair. The circles are qubits, the solid lines represent the entanglement between qubits and the dotted lines represent the extension of the cluster following the presented structure, so only part of a large cluster is explicitly drawn. At the top, the initial situation is depicted, where the target qubits (filled circles) have been chosen. Also, the corresponding manipulation is established, which consists of the measurement in the $\sigma_{z}$ basis of the outer neighbors and applying a measurement pattern to the inner neighbors. At the bottom, the manipulation has been performed and the resulting Bell pair and the remaining graph state are presented.
3.3 One-dimensional cluster as a building block